OpenStudy (anonymous):

can someone help me with this problem? √(5x-1)=√(4x+9)

OpenStudy (anonymous):

\[5x-1=4x+9\] solve for x

OpenStudy (anonymous):

square both sides, so 5x-1=4x+9, then get all the x's to one side, 5x-4x-1=9, then add 1 to each side 5x-4x=10 then subtract 4x from 5x so x=10